OpenStudy (anonymous):

I'm trying to check my answers. Sodium Hydroxide is formed when sodium oxide reacts with water. Na2O + H2O -->2NaOH a) Calculate the grams of sodium hydroxide formed when 2.24 moles of sodium oxide react with water. My answer: 179.2 grams NaOH b) What mass of water (in centigrams) is needed to react with 0.216 mol of sodium oxide? My answer: 468 centigrams H2O

OpenStudy (somy):

First part is right Second part im getting this ratio between Na2O and H2O is 1:1 so H2O's mole will be same =0.216 mol mole= mass / Mr mass= mole*Mr =0.126*18 =3.888 g convert to centigrams= 388.8 centigrams I don't know though, maybe you are right and im wrong :)